OpenStudy (anonymous):

4668.97 + -327.5 + -810.78 + - 269.75 + - 278.55 =

OpenStudy (anonymous):

4668.97 + (-327.5) + (- 810.78) + (-269.75) + (-278.55)= 4668.97 - 327.5 - 810.78 - 269.75 - 278.55 = 4668.97 - 1686.58 = 2982.39